OpenStudy (pants05):

The lcm of the two numbers is 12. The sum of the two numbers is 10. What is the value of the greater of the two numbers?

OpenStudy (sshayer):

factors of 12 are 2,3,4,6 L.C.M. of 3 and 4=12 but 3+4 =7(not 10) L.C.M. of 4 and 6=12 and 4+6=10 so value of greater number=6

OpenStudy (sshayer):

i am leaving now